### v1

# Learning the Influence Graph of a Markov Process that Randomly Resets to Past

Abstract
Learning the influence graph $G$ of a high-dimensional Markov processes is a challenging problem. Prior work [1] addresses this task when the process has finite memory. However, the more general regime in which the system probabilistically “jumps back in time”—so that the state at t+1t{+}1 depends on a sample from a distant past tdt{-}d—remains unexplored. The process with probabilistic resets can be modeled as a Markov process with memory, but estimations become computationally expensive. To tackle this, this letter introduces PIMRecGreedy, a modification of the RecGreedy algorithm from [2], originally designed for i.i.d. samples. The proposed method doesn’t assume memory, therefore requires no prior knowledge of $d$ and recovers $G$ with high probability even without access to the specific time indices at which such temporal jumps occur, and without imposing any constraints on the graph structures.
I Introduction
Learning the underlying influence (causal) graph $G$ of a high-dimensional multivariate stochastic process has applications in multiple domains, including online social networks, neural and biological signal processing, and financial risk analysis [3, 4, 5, 6]. The existing literature focuses on learning the influence graphs of Markov stochastic processes. Motivated by important phenomena in social networks and financial risk analysis, in this paper we study the problem of learning the influence graphs of a class of stochastic processes whose distant past occasionally influences the future in an unpredictable way.
In a multivariate Markov dynamic with an influence graph, the next state of a process variable depends on the current states of process variables that have directed edges to it. We study the following modification of such a Markov process. The future states depend on the current states with probability $p$, and with probability 1p1-p they depend on the states $d$ time-steps back. Intuitively, it is a process in which a Markovian evolution is randomly reset to a distant past state. The task is to recover the influence graph from samples of the process observed over $T$ time-steps, without knowing the instances when the “random resets to the past” occurred.
This phenomenon appears in many domains, where the distant past occasionally influences the future in addition to present. For instance, in social networks, users may sometimes form opinions based on older interactions rather than the latest activities. In many social networks, agents express short opinions in binary form (likes/dislikes), which in turn influence the beliefs of others in the network. Occasionally, opinions expressed in the distant past resurface and shape the opinions of others. Similarly, in financial markets, sudden shocks or crises often resurface after long intervals, influencing present-day risk perception and investment behavior. For example, an unexpected downturn may trigger responses not only to recent market fluctuations but also to memories of past crises (such as the 2008 financial collapse), leading investors to update their portfolios based on distant past rather than current scenario [7, 8].
The social network model proposed in this paper captures the phenomenon in which the network as a whole is influenced by the past, rather than just a subset of participants. We also model the reactions of the network participants with a binary alphabet, though the framework can be easily extended to problems involving larger observable symbol sets.
Though analytical guarantees for learning processes with random resets to past have not been studied before, a large body of work has addressed the problem of learning i.i.d. graphical model structures. Early work by Chow and Liu [9] introduced tree-structured approximations for discrete distributions, and subsequent efforts extended these ideas to bounded tree-width Markov networks [10]. In high-dimensional settings, algorithms based on sparsity have been popular, including $\ell_{1}$-regularized logistic regression for Ising models [11] and group-sparse regularization techniques [12]. Other approaches exploit structural properties: Bresler et al. [13] proposed reconstruction methods for Markov random fields, while Wu et al. [14] investigated loosely connected graphs. Greedy methods have also proven effective for scalable structure recovery [15]. And, more recently [16] provided algorithms for learning directed Gaussian graphical models with sample complexity.
The algorithm proposed in this paper is for learning processes with random resets to past, which is a generalization of the RecGreedy algorithm introduced in [2] for i.i.d. samples. Recently, a similar approach was adopted in [1] to learn Markovian processes without random resets to past, but with a finite memory.
The paper is organized as follows. In Section II we introduce the Past Influence Model, and describe how its evolves over discrete time steps. In Section III we define the learning problem, followed by Section IV, where we propose the PIMRecGreedy algorithm, explaining how it works and why it is more effective than simply considering a Markov process with a large enough memory, along with its sample complexity theorem. Section V presents the performance of PIMRecGreedy across varying sample sizes The proof of Theorem 1 is provided in the Appendix.
II The Social Network Model
II-A Modeling
Influence is modeled by a directed graph $G=(V,E)$, where nodes represent agents and directed edges represent influence. A directed edge $(u,v)\in E$ means $u$ influences $v$, and self-influence is modeled by $(v,v)\in E$. The neighborhood of $v$ is $\mathcal{N}_{v}:=\{u\neq v:(u,v)\in E\}$.
Time evolves in discrete steps. At each $t$, node $v$ generates $M_{v}(t)\sim\min(\text{Poisson}(\mu_{v}(X_{v}(t))),\overline{M})+1$ independent Bernoulli samples with success probability $X_{v}(t)$, where $\mu_{v}:[0,1]\to\mathbb{R}$ is $L$-Lipschitz and $\overline{M}$ is a constant cap. Out of these samples, $N_{v}(t)$ are ones.
Each directed edge $(u,v)$ has weight $a_{uv}\in(0,1)$, with $\sum_{u}a_{uv}=1$. The internal parameter $X_{v}(t)$ evolves via Eqn. (II-A) through three components: neighborhood observations $\mathcal{N}_{v}$, intrinsic bias $l_{v}$, and random fluctuation $Z_{v}(t)\in[0,1]$ with mean $0<\bar{z}_{v}<1$ and parameter $\beta\in(0,1)$. $C(t)\sim\mathrm{Ber}(p)$ models the ignoring of the last $d$ samples. The coefficient $\alpha_{v}$ measures $v$’s openness to external influence, trading off social input against internal dynamics. We call this the PIM (Past Influence Model). We sometimes refer to the time-resets as tails.
$X_{v}(t+1)=(1-\alpha_{v})\left[(1-\beta)Z_{v}(t)+\beta l_{v}\right]$
$+\alpha_{v}\sum_{u\in\mathcal{N_{v}}\cup v}\alpha_{uv}\left[C(t)\frac{N_{u}(t)}{M_{u}(t)}+(1-C(t))\frac{N_{u}(t-d)}{M_{u}(t-d)}\right]$ (1)
II-B Some Mathematical Structures
Given the observations $\{(N_{v}(t),M_{v}(t)):v\in V\}_{t=0}^{T-1}$, we define the variable $Y_{v}(t):=\frac{N_{v}(t)}{M_{v}(t)}$. We also define:
  • $Y_{Q}(t):=\{Y_{v}(t):v\in Q\subseteq V\}$
  • $Y_{u,Q}(t):=\{Y_{v}(t):v\in Q\cup\{u\},Q\subseteq V\}$
  • $Y(t):=\{Y_{v}(t):v\in V\}$
If there was no coin toss in Eqn. (II-A) the process $Y(t)$ becomes a normal Markov chain, $\{W(t)\}$. The directed conditional entropy of node $v$ given a set $Q$ is defined as:
$\sum_{Y_{Q,v}(t)}\sum_{Y_{v}(t+1)}\mathbf{P}(Y_{v}(t+1),Y_{Q,v}(t))\log\mathbf{P}(Y_{v}(t+1)|Y_{Q,v}(t)),$
for some $Q\subset V$ and $v\not\in Q$. In a stationary setting, this quantity becomes time-invariant and is denoted as $H(v_{+}|v,Q)$. This can be computed by first calculating the joint entropies $H(v_{+},v,Q)=-\sum\mathbf{P}(Y_{v}(t+1),Y_{Q,v}(t))\log\mathbf{P}(Y_{v}(t+1),Y_{Q,v}(t))$ and $H(v,Q)$ and $H(v,Q)=-\sum_{Y_{Q,v}(t)}\mathbf{P}(Y_{Q,v}(t))\log\mathbf{P}(Y_{Q,v}(t))$, in the former, the summation is over $Y_{Q,v}(t)$ and $Y_{v}(t+1)$.
In practice, the conditional entropy $H(v_{+}|v,Q)$ cannot be computed directly due to unknown distributions. Instead, we estimate it from data and denote the estimate by $\hat{H}(v_{+}|v,Q)$. Specifically, we compute plug-in estimates of the joint entropies $\hat{H}(v_{+},v,Q)$ and $\hat{H}(v,Q)$ using the empirical distributions $\hat{p}(y_{v},y_{Q})$ and $\hat{p}(y_{v_{+}},y_{v},y_{Q})$, and then take their difference.
II-C Probabilistically teleporting to the past
At each discrete time step $t$, the PIM, according to Eqn. (II-A), performs a Bernoulli trial with probability of success $p$. If the outcome is a tail, the model ignores the most recent $d$ samples and instead evolves based on the samples from time tdt-d. Hence, the model effectively follows the distribution from time td+1t-d+1, which we refer to as a reset to the past. In this way, the true state of $Y(t+1)$ aligns with the true state at time td+1t-d+1.
$y(1),y(2),\ldots,y(d+1),\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{y^{(2)}(2)},\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{y^{(2)}(3)},\ldots,y^{(2)}(d+1),$ (2)
$y(d+2),\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{y^{(3)}(3)},\ldots,\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{y^{(3)}(d+2)},y(d+3),\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{y^{(4)}(4)},\ldots$
In the above run of $Y(t)$, the model forgets the recent past at the $(d+2)^{\text{th}}$ time step. As a result, the sample $Y^{(2)}(2)$ is drawn from the same distribution as $Y(2)$; similarly, $Y^{(3)}(3)$ is distributed identically to both $Y^{(2)}(3)$ and $Y(3)$.
The observations (2) are then processed as in (3) to estimate probabilities. Since the PIM does not know the positions of the tails, mismatched samples (e.g., $y^{(2)}(2)$ generated from $y(1)$) are also erroneously included in the estimation. Nevertheless, with sufficiently large sample size (cf. Theorem 1), the estimation error remains small enough to reliably separate neighbors from non-neighbors.
$[y(1),y(2)],\ldots,\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{[y(d+1),y^{(2)}(2)]},\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{[y^{(2)}(2),y^{(2)}(3)]},\ldots$ (3)
$[y^{(2)}(d+1),y(d+2)],\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{[y(d+2),y^{(3)}(3)]},[y^{(3)}(3),y^{(3)}(4)],\ldots$
$\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{[y^{(2)}(d+1),y^{(3)}(d+2)]},\ldots,\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{[y(d+3),y^{(4)}(4)]},\ldots$
III The Learning Problem
[…]
IV Algorithm to Recover the Influence Graph
In this section, we describe a greedy approach to recover the underlying directed graph $G$ using the observation sequence $\{(N_{v}(t),M_{v}(t)):v\in V\}_{t=0}^{T-1}$. This procedure is a variant of the RecGreedy algorithm proposed for learning graphical models with i.i.d. samples [2]. Our primary technical contribution lies in establishing a sample complexity guarantee in the context of Markovian dynamics that probabilistically forgets the recent samples. Specifically, we demonstrate that the algorithm can recover the true graph with high probability using data collected over $O(\log|V|)$ time steps.
Algorithm 1 may seem similar to the algorithm proposed in [1], but [1] relies on estimating probabilities of the form $\hat{p}(Y_{v+}\mid Y^{m}_{v,Q})$, where
$Y^{m}_{v,Q}(t)=\{Y_{v,Q}(i):i\in[t-m,t]\},$
where $m$ is the Markov chain’s memory. For higher values of $d$, following the approach of [1] with $m=d$ becomes computationally very challenging, as probability estimation is required over large conditioning sets. By contrast, the PIMRecGreedy algorithm presented here assumes no memory. When the sample size $T$ satisfies Theorem 1, the algorithm recovers the graph with very high probability.
The algorithm relies on estimates of directed conditional entropy, $\hat{H}(v_{+}\mid v,Q)$, for various $Q\subset V$. For each node $v$, it begins with empty sets $\hat{U}(v)$ and $\hat{T}(v)$, representing intermediate and final neighborhood estimates. At each iteration, the algorithm selects the node $u$ that maximizes the reduction in directed conditional entropy, i.e.,
$\hat{H}(v_{+}\mid v,\hat{U}(v))-\hat{H}(v_{+}\mid v,\hat{U}(v),u),$
and includes $u$ in $\hat{U}(v)$ if the reduction exceeds a fixed threshold. This process continues until no further node yields a reduction exceeding the threshold.
When the threshold is set to $\kappa/2$, Lemma 2 guarantees that the last node added to $\hat{U}(v)$ is always a true neighbor of $v$. Only this last added node is included in the final estimate $\hat{T}(v)$. The procedure is then repeated, initializing with $\hat{U}(v)=\hat{T}(v)$. The process terminates once $\hat{T}(v)$ matches the true neighborhood, as no additional nodes satisfy the threshold condition.
The algorithm’s sample complexity is largely determined by the dimension of the probability distributions that need to be estimated from data. Since it proceeds from smaller subsets (P) and stops when no further decrease in entropy is observed, the dimension of the involved distributions remains bounded.
To analyze the sample complexity, we introduce
$|P|_{max}\leq\frac{2\log\!\left(\tfrac{\bar{M}\left(\bar{M}+1\right)}{2}+2\right)}{\epsilon^{\prime}}+1$
(Lemma 5, [1]) [^1] The bound is obtained by exploiting the fact that the inner loop of the algorithm reduces the entropy by at least $\epsilon^{\prime}/2$ upon adding a new node to $\hat{U}(v)$. to denote the largest cardinality of subset $P$ required. Let $\chi$ denote the support of the random variable $Y_{v}(t)$, and define $|\xi|:=|\chi|^{(2+|P|_{max})}$ to be the maximum support size of the joint distributions to be estimated. Together with properties of the Markov process, this yields a logarithmic sample complexity, formally established in Theorem 1.
Let $p_{g}$ denote the probability distribution of the observed samples given the instants of resets known to the observer and the observations are ordered (in time) accordingly, we shall call this the post-processing distribution. This distribution is used only for the proof of sample complexity and won’t be needed for the implementation. For $M_{v}(t)=1$. it can be shown that $\sum_{w_{v_{+}}}|p_{g}(y_{v_{+}}|y_{v,Q})-p_{g}(y_{v_{+}}|y_{v,Q,u})|$ is bounded away from zero, for $\forall v\in V$, any values of $y_{v},y_{Q}$, when $u\in\mathcal{N}_{v}$, and $Q\subset V$ such that $u\not\in Q$. For the general case, due to certain technicalities, we make the following assumption.
Assumption 1.
Consider any $v\in V$, $u\in\mathcal{N}_{v}$, and $Q\subset V$ such that $u\not\in Q$. Then for any values $y_{v}$ and $y_{Q}$, there exists a value $y_{u}$ such that for some constant $\kappa^{\prime}>0$,
$\sum_{y_{v_{+}}}|p_{g}(y_{v_{+}}|y_{v,Q})-p_{g}(y_{v_{+}}|y_{v,Q,u})|>\sqrt{\kappa^{\prime}}$
Theorem 1.
[…]
$T\geq\max\Biggl\{\log\!\left(\frac{c+2\cdot|V|^{(P_{\max}+1)}\,|\xi|}{\gamma}\right)\frac{12\delta^{\prime}\beta_{1}}{c_{1}^{2}},\;$
$(d^{2}-d+2)+\frac{(1-2(\bar{\mu}+L)|\rho(\tilde{A})|)\delta^{2}}{2(1+2(\bar{\mu}+L)|\rho(\tilde{A})|)|\xi|^{2}(d+1)^{-1}}$
$\quad\times\log\!\left(\frac{c+2\cdot|V|^{(P_{\max}+1)}\,|\xi|}{\gamma}\right)\Biggr\}$
Where, $c>0,|\xi|:=|\chi|^{(2+|P|_{max})}$, $|\chi|\leq(\frac{\bar{M}(\bar{M}+1)}{2}+2)$, $\beta_{1}\delta^{\prime}-4(T-1)^{\alpha-1}>c_{1}$ $\delta.\log\left(\frac{|\xi|}{\delta}\right)\leq\frac{\epsilon}{4}$, $\delta^{\prime}.\log\frac{|\xi|}{\delta^{\prime}}\leq\frac{\epsilon^{\prime}}{4}$ and $(1-p)=\frac{(T-1)^{\alpha}}{\beta_{1}(T-d-1)}$, for some $\alpha<1,\text{ and }0<\beta_{1}<1$, $2(\bar{\mu}+L)|\rho(\tilde{A})|<1$, and Assumption 1 holds.
Algorithm 1 PIMRecGreedy($\kappa$) Algorithm
Input: $\{N_{v}(t):v\in V\}_{t=0}^{T-1},\{M_{v}(t):v\in V\}_{t=0}^{T-1},$
   Output: $\hat{T}(v):\forall v\in V$}
1:for $v\in V$ do
2:  $\hat{T}(v),\leftarrow\phi$
3:  repeat
4:   $\hat{U}(v)\leftarrow\hat{T}(v)$
5:   $\text{LastNode}\leftarrow\phi$
6:   repeat
7:
$u$ $=\arg\max\limits_{k\in V\setminus\hat{U}(v)}\hat{H}(v_{+}|v,\hat{U}(v))$
$-\hat{H}(v_{+}|v,\hat{U}(v),k)$
8:     $\delta^{\prime}_{u}=\hat{H}(v_{+}|v,\hat{U}(v))-\hat{H}(v_{+}|v,\hat{U}(v),u)$
9:     if $\delta^{\prime}_{u}>\kappa/2$ then
10:      $\hat{U}(v)\leftarrow\hat{U}(v)\cup\{u\}$
11:      $\text{LastNode}\leftarrow u$
12:     else
13:      $\hat{T}(v)\leftarrow\hat{T}(v)\cup\text{LastNode}$
14:     end if
15:   until New node is not added to $\hat{U}(v)$
16:  until New node is not added to $\hat{T}(v)$
17:end for
V Simulation results
The data are generated using Eqn. (II-A) with parameters $l_{v}=0.167$, $\beta=0.75$, $\alpha_{v}=0.8$, $\beta_{1}=0.75$, $\alpha=0.5$, $|V|=10$, $\mu_{v}(x)=0.4x$, and $\alpha_{uv}=1/|\mathcal{N}_{v}|$ for $u\in\mathcal{N}_{v}$. Performance of PIMRecGreedy over ring and line graphs are plotted in Fig. 1 for $d=5$ and $\overline{M}=1$. The simulations are repeated for a slightly harder case with $d=10$ and $\overline{M}=2$, and the results are shown in Fig. 2.
Choice of Parameter $\kappa$
The threshold parameter $\kappa$ for PIMRecGreedy is selected via cross-validation on a labeled/held-out dataset, by choosing the value that maximizes a suitable performance metric (here, the probability of full graph recovery). Similar parameter tuning is required in other graphical model learning methods, such as the regularization parameter $\lambda$ in $RWL$ [11] and the local separator size $\eta$ in the $CVDT$ algorithm [17]. Cross-validation results on ring and line graphs with 10 nodes are shown in Fig.3 and Fig.4, corresponding to $d=5,\overline{M}=1$ and $d=10,\overline{M}=2$, respectively
VI Conclusion
We have introduced a recursive greedy algorithm, PIMRecGreedy, that can recover the influence graph of a high-dimensional Markov process which, at every time step, probabilistically forgets its recent samples. PIMRecGreedy is robust in the sense that it does not require prior knowledge of how far into the past the process travels, nor does it assume information about where temporal jumps occur in the given samples. This makes the algorithm significantly more computationally feasible than implementing a greedy algorithm for a Markov process with memory.
Figure 1: $d=5$ and $\overline{M}=1$
Figure 2: $d=10$ and $\overline{M}=2$
Figure 3: $d=5$,$T=3000$, and $\overline{M}=1$
Figure 4: $d=10$, $T=4000$, and $\overline{M}=2$

### v2

@@
-Learning the Influence Graph of a Markov Process that Randomly Resets to Past
+Learning the Influence Graph of a Markov Process that Randomly Resets to the Past
 Abstract
-Learning the influence graph $G$ of a high-dimensional Markov processes is a challenging problem.
-Prior work [1] addresses this task when the process has finite memory.
-However, the more general regime in which the system probabilistically “jumps back in time”—so that the state at t+1t{+}1 depends on a sample from a distant past tdt{-}d—remains unexplored.
-The process with probabilistic resets can be modeled as a Markov process with memory, but estimations become computationally expensive.
-To tackle this, this letter introduces PIMRecGreedy, a modification of the RecGreedy algorithm from [2], originally designed for i.i.d.
-samples.
-The proposed method doesn’t assume memory, therefore requires no prior knowledge of $d$ and recovers $G$ with high probability even without access to the specific time indices at which such temporal jumps occur, and without imposing any constraints on the graph structures.
+Learning the influence graph $G$ of a high-dimensional Markov process is central to many application domains, including social networks, neuroscience, and financial risk analysis.
+However, in many of these applications, future states of the process are occasionally and unpredictably influenced by a distant past state, thus destroying the Markovianity.
+To study this practical issue, we propose the past influence model (PIM), which captures the occasional ”random resets to past” by modifying the Markovian dynamics in [2], which, in turn, is a non-linear generalization of the dynamics studied in [13], [4].
+The recursive greedy algorithm proposed in this paper recovers any bounded degree $G$ when the number of “jumps back in time” is order-wise smaller than the total number of samples, and the algorithm does not require memory.
 I Introduction
@@
 The existing literature focuses on learning the influence graphs of Markov stochastic processes.
-Motivated by important phenomena in social networks and financial risk analysis, in this paper we study the problem of learning the influence graphs of a class of stochastic processes whose distant past occasionally influences the future in an unpredictable way.
+Motivated by important phenomena in social networks and financial risk analysis, this paper studies the problem of learning the influence graphs of a class of stochastic processes whose distant past unpredictably influences their future.
 In a multivariate Markov dynamic with an influence graph, the next state of a process variable depends on the current states of process variables that have directed edges to it.
@@
 The task is to recover the influence graph from samples of the process observed over $T$ time-steps, without knowing the instances when the “random resets to the past” occurred.
-This phenomenon appears in many domains, where the distant past occasionally influences the future in addition to present.
+This phenomenon is observed in many domains, where the distant past occasionally influences the future, in addition to the present.
 For instance, in social networks, users may sometimes form opinions based on older interactions rather than the latest activities.
-In many social networks, agents express short opinions in binary form (likes/dislikes), which in turn influence the beliefs of others in the network.
+In many social networks, agents express their opinions in a binary form (likes/dislikes), which in turn influences the beliefs of others within the network.
 Occasionally, opinions expressed in the distant past resurface and shape the opinions of others.
 Similarly, in financial markets, sudden shocks or crises often resurface after long intervals, influencing present-day risk perception and investment behavior.
-For example, an unexpected downturn may trigger responses not only to recent market fluctuations but also to memories of past crises (such as the 2008 financial collapse), leading investors to update their portfolios based on distant past rather than current scenario [7, 8].
-The social network model proposed in this paper captures the phenomenon in which the network as a whole is influenced by the past, rather than just a subset of participants.
+The social network model proposed in this paper addresses the situation in which the entire network is influenced by its past, rather than just a subset of participants.
 We also model the reactions of the network participants with a binary alphabet, though the framework can be easily extended to problems involving larger observable symbol sets.
-Though analytical guarantees for learning processes with random resets to past have not been studied before, a large body of work has addressed the problem of learning i.i.d.
+Though analytical guarantees for learning processes with random resets to the past have not been studied before to our best knowledge, a large body of work has addressed the problem of learning i.i.d.
 graphical model structures.
@@
 And, more recently [7] provided algorithms for learning directed Gaussian graphical models with sample complexity.
-The algorithm proposed in this paper is for learning processes with random resets to past, which is a generalization of the RecGreedy algorithm introduced in [2] for i.i.d.
+The algorithm proposed in this paper is for learning processes with random resets to the past, which is a generalization of the RecGreedy algorithm introduced in [18] for i.i.d.
 samples.
-Recently, a similar approach was adopted in [1] to learn Markovian processes without random resets to past, but with a finite memory.
-The paper is organized as follows.
-In Section II we introduce the Past Influence Model, and describe how its evolves over discrete time steps.
+Recently, a similar approach was adopted in [2] to learn Markovian processes without random resets to the past, but with a finite memory.
+The paper is organized as follows: In Section II we introduce the Past Influence Model, and describe how it evolves over discrete time steps.
 In Section III we define the learning problem, followed by Section IV, where we propose the PIMRecGreedy algorithm, explaining how it works and why it is more effective than simply considering a Markov process with a large enough memory, along with its sample complexity theorem.
-Section V presents the performance of PIMRecGreedy across varying sample sizes The proof of Theorem 1 is provided in the Appendix.
+Section V presents the performance of PIMRecGreedy across varying sample sizes.
 II The Social Network Model
@@
 (II-A) through three components: neighborhood observations $\mathcal{N}_{v}$, intrinsic bias $l_{v}$, and random fluctuation $Z_{v}(t)\in[0,1]$ with mean $0<\bar{z}_{v}<1$ and parameter $\beta\in(0,1)$.
-$C(t)\sim\mathrm{Ber}(p)$ models the ignoring of the last $d$ samples.
+$C(t)\sim\mathrm{Ber}(p)$ models the random reset to the past (td)(t-d).
 The coefficient $\alpha_{v}$ measures $v$’s openness to external influence, trading off social input against internal dynamics.
 We call this the PIM (Past Influence Model).
-We sometimes refer to the time-resets as tails.
+We sometimes refer to the time-reset as occurrence of Tail (in a coin toss perspective).
 $X_{v}(t+1)=(1-\alpha_{v})\left[(1-\beta)Z_{v}(t)+\beta l_{v}\right]$
@@
 In this way, the true state of $Y(t+1)$ aligns with the true state at time td+1t-d+1.
-$y(1),y(2),\ldots,y(d+1),\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{y^{(2)}(2)},\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{y^{(2)}(3)},\ldots,y^{(2)}(d+1),$ (2)
-$y(d+2),\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{y^{(3)}(3)},\ldots,\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{y^{(3)}(d+2)},y(d+3),\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{y^{(4)}(4)},\ldots$
+$y(1),y(2),..,y(d+1),\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{y^{(2)}(2)},\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{y^{(2)}(3)},..,y^{(2)}(d+1),$ (2)
+$y(d+2),\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{y^{(3)}(3)},..,\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{y^{(3)}(d+2)},y(d+3),\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{y^{(4)}(4)},..$
 In the above run of $Y(t)$, the model forgets the recent past at the $(d+2)^{\text{th}}$ time step.
@@
 Theorem 1), the estimation error remains small enough to reliably separate neighbors from non-neighbors.
-$[y(1),y(2)],\ldots,\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{[y(d+1),y^{(2)}(2)]},\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{[y^{(2)}(2),y^{(2)}(3)]},\ldots$ (3)
-$[y^{(2)}(d+1),y(d+2)],\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{[y(d+2),y^{(3)}(3)]},[y^{(3)}(3),y^{(3)}(4)],\ldots$
-$\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{[y^{(2)}(d+1),y^{(3)}(d+2)]},\ldots,\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{[y(d+3),y^{(4)}(4)]},\ldots$
+$[y(1),y(2)],..,\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{[y(d+1),y^{(2)}(2)]},\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{[y^{(2)}(2),y^{(2)}(3)]},..$ (3)
+$[y^{(2)}(d+1),y(d+2)],\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{[y(d+2),y^{(3)}(3)]},[y^{(3)}(3),y^{(3)}(4)],..$
+$\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{[y^{(2)}(d+1),y^{(3)}(d+2)]},..,\underset{\textbf{{\color[rgb]{1,0,0}Tail}}}{[y(d+3),y^{(4)}(4)]},..$
 III The Learning Problem
@@
 samples [18].
-Our primary technical contribution lies in establishing a sample complexity guarantee in the context of Markovian dynamics that probabilistically forgets the recent samples.
-Specifically, we demonstrate that the algorithm can recover the true graph with high probability using data collected over $O(\log|V|)$ time steps.
-Algorithm 1 may seem similar to the algorithm proposed in [1], but [1] relies on estimating probabilities of the form $\hat{p}(Y_{v+}\mid Y^{m}_{v,Q})$, where
+Our primary technical contribution lies in establishing a sample complexity guarantee for Markovian dynamics that probabilistically gets reset to the past.
+Specifically, we demonstrate that the algorithm can recover the true graph with high probability using data collected over $\mathcal{O}(d^{2}+d\log|V|)$ time steps.
+Algorithm 1 is similar in spirit to the algorithm proposed in [2], but that relies on estimating entropies of the form $\hat{H}(Y_{v+}\mid Y^{m}_{v,Q})$, where
 $Y^{m}_{v,Q}(t)=\{Y_{v,Q}(i):i\in[t-m,t]\},$
@@
 For higher values of $d$, following the approach of [2] with $m=d$ becomes computationally very challenging, as probability estimation is required over large conditioning sets.
-By contrast, the PIMRecGreedy algorithm presented here assumes no memory.
-When the sample size $T$ satisfies Theorem 1, the algorithm recovers the graph with very high probability.
+By contrast, the PIMRecGreedy algorithm presented here requires no memory.
+When the sample size $T$ satisfies Theorem 1, the algorithm recovers the graph with a very high probability.
 The algorithm relies on estimates of directed conditional entropy, $\hat{H}(v_{+}\mid v,Q)$, for various $Q\subset V$.
@@
 The algorithm’s sample complexity is largely determined by the dimension of the probability distributions that need to be estimated from data.
-Since it proceeds from smaller subsets (P) and stops when no further decrease in entropy is observed, the dimension of the involved distributions remains bounded.
-To analyze the sample complexity, we introduce
-$|P|_{max}\leq\frac{2\log\!\left(\tfrac{\bar{M}\left(\bar{M}+1\right)}{2}+2\right)}{\epsilon^{\prime}}+1$
-(Lemma 5, [1]) [^1] The bound is obtained by exploiting the fact that the inner loop of the algorithm reduces the entropy by at least $\epsilon^{\prime}/2$ upon adding a new node to $\hat{U}(v)$.
-to denote the largest cardinality of subset $P$ required.
+Since it proceeds from smaller subsets P and stops when no further decrease in entropy is observed, the dimension of the involved distributions remains bounded.
+To analyze the sample complexity, we use (Lemma 5, [2]),
+$|P|_{max}\leq\frac{2\log\!\left(\tfrac{\bar{M}\left(\bar{M}+1\right)}{2}+2\right)}{\kappa}+1$
+to denote the largest cardinality of subset $P$ required  [^1] The bound is obtained by exploiting the fact that the inner loop of the algorithm reduces the entropy by at least $\kappa/2$ upon adding a new node to $\hat{U}(v)$..
 Let $\chi$ denote the support of the random variable $Y_{v}(t)$, and define $|\xi|:=|\chi|^{(2+|P|_{max})}$ to be the maximum support size of the joint distributions to be estimated.
 Together with properties of the Markov process, this yields a logarithmic sample complexity, formally established in Theorem 1.
-Let $p_{g}$ denote the probability distribution of the observed samples given the instants of resets known to the observer and the observations are ordered (in time) accordingly, we shall call this the post-processing distribution.
-This distribution is used only for the proof of sample complexity and won’t be needed for the implementation.
+Let $p_{g}$ denote the probability distribution of the observed samples given the instants of tails and the observations are ordered (in time) accordingly, we shall call this the post-processing distribution.
+This distribution is used only for the sample complexity analysis.
 For $M_{v}(t)=1$.
@@
 The data are generated using Eqn.
-(II-A) with parameters $l_{v}=0.167$, $\beta=0.75$, $\alpha_{v}=0.8$, $\beta_{1}=0.75$, $\alpha=0.5$, $|V|=10$, $\mu_{v}(x)=0.4x$, and $\alpha_{uv}=1/|\mathcal{N}_{v}|$ for $u\in\mathcal{N}_{v}$.
+(II-A) with parameters $l_{v}=0.167$, $\beta=0.75$, $\alpha_{v}=0.8$, $\beta_{1}=0.75$, $\alpha=0.5$, $|V|=5\text{ and }10$, $\mu_{v}(x)=0.4x$, and $\alpha_{uv}=1/|\mathcal{N}_{v}|$ for $u\in\mathcal{N}_{v}$.
 Performance of PIMRecGreedy over ring and line graphs are plotted in Fig.
@@
 2.
+The simulation results for the graphs presented in Fig.
+3 are shown in Fig.
+4.
 Choice of Parameter $\kappa$
@@
 Similar parameter tuning is required in other graphical model learning methods, such as the regularization parameter $\lambda$ in $RWL$ [17] and the local separator size $\eta$ in the $CVDT$ algorithm [1].
-Cross-validation results on ring and line graphs with 10 nodes are shown in Fig.3 and Fig.4, corresponding to $d=5,\overline{M}=1$ and $d=10,\overline{M}=2$, respectively
+Cross-validation results on ring and line graphs with 10 nodes are shown in Fig.
+5, for $d=10\text{ and }\overline{M}=2$.
+Figure 1: $|V|=5,d=5$, and $\overline{M}=1$
+Figure 2: $|V|=10,d=10$, and $\overline{M}=2$
+Figure 3: A) $K_{3}$ with a pendant, B) $S_{3}$, and C) Tree $|V|=5$.
+Figure 4: $d=10$ and $\overline{M}=1$
+Figure 5: $|V|=10,d=10$, $T=3500$, and $\overline{M}=2$
 VI Conclusion
-We have introduced a recursive greedy algorithm, PIMRecGreedy, that can recover the influence graph of a high-dimensional Markov process which, at every time step, probabilistically forgets its recent samples.
-PIMRecGreedy is robust in the sense that it does not require prior knowledge of how far into the past the process travels, nor does it assume information about where temporal jumps occur in the given samples.
-This makes the algorithm significantly more computationally feasible than implementing a greedy algorithm for a Markov process with memory.
-Figure 1: $d=5$ and $\overline{M}=1$
-Figure 2: $d=10$ and $\overline{M}=2$
-Figure 3: $d=5$,$T=3000$, and $\overline{M}=1$
-Figure 4: $d=10$, $T=4000$, and $\overline{M}=2$
+In many applications, such as social networks and financial risk analysis, a distant past state can occasionally and randomly affect the future.
+Motivated by this, we propose the past influence model (PIM) by adapting a general Markovian model [2, 13] and propose an algorithm to learn the underlying influence graph of this process with a provable guarantee on accuracy and sample complexity.
+The algorithm has a certain threshold $\kappa$ (similar to regularization parameters), and we observe that, in experiments (Fig.
+5), the algorithm’s performance depends on this threshold.
+It was also observed that the influence graph could be learned for constant values of 1p1-p (Probability of resetting), but its analysis is deferred for future work.
+This work is just the first step towards generalizing this model to capture the specific issues of social networks more effectively.
+It would be a challenging problem, but with direct applications.